Please do all the parts and explain it well

Priya has Cobb Douglas preferences for Juice (X1) and Soda (X2). Her MRS at his optimal consumption is 2. If the price of juice is p1= $6. What is the price of Soda, p2?
Answer #1

At the optimal consumption,

MRS = price ratio ( Price of juice / Price of Soda)

Put values, we get

MRS = p1/ p2

2 = 6/ p2

p2 = 6/2 = 3

So price of Soda, P2 = $3.